OpenStudy (anonymous):

Solve the system x=7y-4 y=7x+4 for X

OpenStudy (dumbcow):

substitution: x = 7(7x+4) -4

OpenStudy (anonymous):

2/3

OpenStudy (dumbcow):

just giving a number answer doesn't help them...you could have just made it up try to show some steps for their sake

OpenStudy (dumbcow):

in fact x = 49x+28 -4 -48x = 24 x = -1/2

OpenStudy (dumbcow):

do you need to solve for y as well?

OpenStudy (anonymous):

Nope, thank you! (: